# Neutrino-nucleon elastic scattering in presence of non-standard interactions: cross sections and nucleon polarizations

### v1

@@ II.2 Neutrino-nucleon NC interactions in presence of NSI
which depend upon the four-momentum transfer squared $Q^{2}=-(k-k^{\prime})^{2}$. Lepton flavor indices have been dropped to alleviate the notation. In the observable quantities, the contribution of $\tilde{F}_{P}$ is always proportional to powers of the outgoing neutrino mass and, therefore, negligible. It is actually exactly zero in the massless neutrino approximation adopted here and disregarded from now on. Owing to isospin symmetry, the FF introduced in Eq. (10) can be related to the isovector and isoscalar vector and axial vector FF. Furthermore, the vector ones can be written in terms of the electromagnetic FF of protons and neutrons $F_{1,2}^{(p,n)}$. In the presence of NSI, one obtains that
@@
$+\left(-\frac{1}{2}\delta_{ij}+\epsilon^{sV}_{ij}\right)F_{1,2}^{s}\,,$
@@ II.3 Form factors
A priori, no conclusive bounds are available from experiments for any of these parameters [59]. On the other hand, significant progress has been made in the determination of nucleon properties on the lattice in the past few years (see for instance the FLAG review [61]). We adopt the numerical values of these parameters obtained in Ref [62] by the Extended Twisted Mass (ETM) Collaboration: $\mu_{s}=-0.017(4)$, $\langle r_{M}^{2}\rangle^{s}=-0.015(9)$ fm2 and $\langle r_{E}^{2}\rangle^{s}=0.0048(6)$ fm2.
@@ III Results
Figure 2: The vector and axial NC form factors defined in Eq. (II.2) are plotted as a function of $Q^{2}$. Solid lines correspond to the SM. The grey bands limited by dashed lines from below and dash-dotted lines from above display the NSI deviations within the 90% CL intervals of Table 2. The green (red) line represents results with (without) strangeness.
The NC vector and axial FF, $\tilde{F}_{1,2,A}^{p,n}$ introduced in Eq. (II.2) are shown in Fig. 2. The solid green (red) line represents the SM results with (without) strangeness. The shaded regions depict the possible deviations due to NSI within the 90% CL intervals of Table 2. It is apparent that, according to the ETM determination of Refs. [62, 63], the impact of strangeness is small compared to the NSI band. Because of the accidental cancellation among terms in the $(1/2-2s_{w}^{2})$ factor, $\tilde{F}_{1}^{p}$ is small in the SM and the possible NSI contribution is relatively large.
Figure 3: Total cross section for NC (anti)neutrino elastic scattering as a function of $E_{\nu}$. The black solid lines represent the SM results while the shaded bands show the deviations by NSI within the 90% CL intervals in Table 2.
@@
In the case of longitudinal ($\zeta_{L}$) polarization it is noteworthy that distinct behaviors are observed for both neutrinos and antineutrinos. The variation due to NSI is nearly negligible close to both low and high $Q^{2}$ thresholds, but a significant variation is evident at intermediate values of $Q^{2}$, particularly for antineutrino scattering. At low $Q^{2}$, the outgoing nucleon’s longitudinal polarization is almost zero and, as $Q^{2}$ increases, it approaches -1 (1) for the neutrino (antineutrino) channel. [^5] We note the different sign of our SM results for the nucleon longitudinal polarization compared to the ones of Ref. [55] (for $\bar{\nu}$ scattering). It is worth stressing that at the largest $Q^{2}$, corresponding to backward scattering, the value of $\zeta_{L}$ is constrained to be -1 (1) for $\nu(\bar{\nu})$ by the conservation of the spin projection along the direction of motion.
On the other hand, in the SM, the transverse component of polarization of the outgoing proton ($\zeta_{T}$) is small at $Q^{2}\sim 0$ and goes to zero quite rapidly. We find that NSI bring significant deviations, specially at low $Q^{2}$. Meanwhile , neutrons are found to be largely polarized in the transverse direction at low $Q^{2}$, while this transverse component gradually reaches zero with increasing $Q^{2}$. Similar findings were observed previously in Ref. [55]. The NSI band appears broader for antineutrinos at middle $Q^{2}$ ranges.
@@ IV Summary and Conclusions
We have studied (anti)neutrino-nucleon NC elastic scattering in presence of NSI of SM neutrinos with light ($u$, $d$, $s$) quarks. Assuming a heavy mediator, these interactions take the well-known form of four-fermion vertices, where the new physics is concealed in the NSI couplings. We have therefore derived the NC matrix elements of the nucleon with NSI, writing them in terms of vector and axial FF, which encode the strong-interaction physics. We find that in presence of NSI, these matrix elements depend not only on the axial isovector FF but also on the isoscalar one, absent in the SM weak interactions. For this reason, for our numerical predictions we rely on the flavor decomposition of the axial FF published by the ETM LQCD collaboration [63]. Under the assumption that there is no significant new physics in electron-nucleon scattering, vector FF can be safely obtained from electron scattering (corrected by radiative corrections). We adopt this strategy for the vector isovector and isoscalar FF but rely on the ETM determination [62] of the strange electromagnetic FF.
In our evaluation of scattering observables, only the muon flavor, with flavor-diagonal, $T$-conserving interactions have been considered. In this scenario there are four non-zero real couplings $\epsilon^{uV,dV}_{\mu\mu}$ and $\epsilon^{uA,dA}_{\mu\mu}$, for which confidence intervals had been obtained from oscillation and deep-inelastic scattering measurements [22]. For such ranges, we find a significant presence of NSI in differential and integrated NC (anti)neutrino-nucleon cross sections, particularly for a proton target and at low $Q^{2}$. For this choice of target and kinematics, NSI have a large impact on the polarization of the outgoing nucleons, especially in the transverse direction.

### v2

@@ II.2 Neutrino-nucleon NC interactions in presence of NSI
which depend upon the four-momentum transfer squared $Q^{2}=-(k-k^{\prime})^{2}$. Lepton flavor indices have been dropped to alleviate the notation. In the observable quantities, the contribution of $\tilde{F}_{P}$ is always proportional to powers of the outgoing neutrino mass and, therefore, negligible. It is actually exactly zero in the massless neutrino approximation adopted here and disregarded from now on. Owing to isospin symmetry, the FF introduced in Eq. (10) can be related to the isovector and isoscalar vector and axial vector FF. Furthermore, the vector ones can be written in terms of the electromagnetic FF of protons and neutrons $F_{1,2}^{(p,n)}$. In the presence of NSI, one obtains that [^2] Although with a different notation, this equation coincides with Eqs. (24-27) of the preprint [59], which appeared in the arXiv a few days after ours.
@@
$+\left(-\frac{1}{2}\delta_{ij}+\epsilon^{uV}_{ij}+\epsilon^{dV}_{ij}+\epsilon^{sV}_{ij}\right)F_{1,2}^{s}\,,$
@@ II.3 Form factors
A priori, no conclusive bounds are available from experiments for any of these parameters [60]. On the other hand, significant progress has been made in the determination of nucleon properties on the lattice in the past few years (see for instance the FLAG review [62]). We adopt the numerical values of these parameters obtained in Ref [63] by the Extended Twisted Mass (ETM) Collaboration: $\mu_{s}=-0.017(4)$, $\langle r_{M}^{2}\rangle^{s}=-0.015(9)$ fm2 and $\langle r_{E}^{2}\rangle^{s}=-0.0048(6)$ fm2.
@@
In Ref. [64], more realistic z-expansion fits were also performed, although their parameters are not disclosed in the publication. Furthermore, the results of the z-expansion fits largely overlap with the dipole ones for $u$ and $d$ quarks, as can be seen in Fig. 15 of Ref. [64]. For these reasons, we rely on the dipole parametrizations although the search for NSI in (future) data will benefit from the best possible input from LQCD.
@@
The errors in the parameters introduced in this section are used to assess the impact of LQCD uncertainties on the observables. We consider these errors to be uncorrelated. Therefore, the resulting bands should be taken with caution as they do not incorporate the correlations that are likely to exist among the LQCD FF.
@@ III Results
Figure 2: The vector and axial NC form factors defined in Eq. (II.2) are plotted as a function of $Q^{2}$. The black solid lines, representing the SM prediction, are accompanied by 90% CL uncertainty bands. The red innermost ones correspond to the errors in the LQCD determination of the axial and strange FF by the ETM Collaboration. The green bands show the deviations by NSI within the 90% CL intervals in Table 2. The outer blue bands arise from the combination of both LQCD and NSI uncertainties.
The NC vector and axial FF, $\tilde{F}_{1,2,A}^{p,n}$ introduced in Eq. (II.2) are shown in Fig. 2. Uncertainties in the LQCD input for the strange vector and flavor-decomposed axial FF, based on the ETM determination of Refs. [63, 64], together with possible deviations due to NSI, have been displayed in the form of 90 % CL bands. It is apparent that the combined uncertainty is dominated by NSI, except at rather high $Q^{2}$. In the vector FF, the small LQCD uncertainty arises solely from strangeness. The contribution of the latter to $\tilde{F}_{A}^{p,n}$ is also very small, so that the LQCD band width is almost entirely driven by the uncertainty in the isovector axial FF. Because of the accidental cancellation among terms in the $(1/2-2s_{w}^{2})$ factor, $\tilde{F}_{1}^{p}$ is small in the SM and the possible NSI contribution is relatively large. We should nevertheless warn the reader about the different abscissa scales in Fig. 2: the NSI band widths are of the same order of magnitude in all the plots.
Figure 3: Total cross section for NC (anti)neutrino elastic scattering as a function of $E_{\nu}$. The black solid line corresponds to the SM. The 90% CL uncertainty bands from LQCD, NSI, and the combination of both, are shown in red, green and blue, respectively.
@@
Next, we compute the total (integrated) cross-sections for the processes given in Eq. (5) to investigate the effects of NSI. The results are shown in Fig. 3. The black solid lines are obtained in the SM, including both vector and axial strange FF, while the red, green and blue-shaded regions give the possible deviation due to LQCD, NSI and LQCD plus NSI, respectively. The cross sections are displayed for incident muon neutrinos and antineutrinos, as well as for protons and neutron targets. The latter would only be experimentally accessible with deuteron targets. In this case, the presence of an additional nucleon would introduce moderate but QCD dependent corrections [69, 70].
@@
In the case of longitudinal ($\zeta_{L}$) polarization it is noteworthy that distinct behaviors are observed for both neutrinos and antineutrinos. The variation due to LQCD or NSI are nearly negligible close to both low and high $Q^{2}$ thresholds, but a significant variation is evident at intermediate values of $Q^{2}$, particularly for antineutrino scattering. The proton polarization in $\bar{\nu}p$ scattering around $Q^{2}=0.2$ GeV displays a relatively large sensitivity to variations in the NSI couplings compared to the LQCD uncertainty. At low $Q^{2}$, the outgoing nucleon’s longitudinal polarization is almost zero and, as $Q^{2}$ increases, it approaches -1 (1) for the neutrino (antineutrino) channel. [^6] We note the different sign of our SM results for the nucleon longitudinal polarization compared to the ones of Ref. [55] (for $\bar{\nu}$ scattering). It is worth stressing that at the largest $Q^{2}$, corresponding to backward scattering, the value of $\zeta_{L}$ is constrained to be -1 (1) for $\nu(\bar{\nu})$ by the conservation of the spin projection along the direction of motion.
On the other hand, in the SM, the transverse component of polarization of the outgoing proton ($\zeta_{T}$) is small at $Q^{2}\sim 0$ and goes to zero quite rapidly. Likewise, the LQCD error is almost negligible. In contrast, we find that NSI bring significant deviations, specially at low $Q^{2}$. Meanwhile, neutrons are found to be largely polarized in the transverse direction at low $Q^{2}$, while this transverse component gradually reaches zero with increasing $Q^{2}$. Similar findings were observed previously in Ref. [55]. Both the LQCD and NSI bands appear broader for antineutrinos at middle $Q^{2}$ ranges.
@@ IV Summary and Conclusions
We have studied (anti)neutrino-nucleon NC elastic scattering in presence of NSI of SM neutrinos with light ($u$, $d$, $s$) quarks. Assuming a heavy mediator, these interactions take the well-known form of four-fermion vertices, where the new physics is concealed in the NSI couplings. We have therefore derived the NC matrix elements of the nucleon with NSI, writing them in terms of vector and axial FF, which encode the strong-interaction physics. We find that in presence of NSI, these matrix elements depend not only on the axial isovector FF but also on the isoscalar one, absent in the SM weak interactions. For this reason, for our numerical predictions we rely on the flavor decomposition of the axial FF published by the ETM LQCD collaboration [64]. The availability of alternative determinations of flavor-diagonal axial form factors would be important for a better assessment of QCD uncertainties. Under the assumption that there is no significant new physics in electron-nucleon scattering, vector FF can be safely obtained from electron scattering (corrected by radiative corrections). We adopt this strategy for the vector isovector and isoscalar FF but rely on the ETM determination [63] of the strange electromagnetic FF.
In our evaluation of scattering observables, only the muon flavor, with flavor-diagonal, $T$-conserving interactions have been considered. In this scenario there are four non-zero real couplings $\epsilon^{uV,dV}_{\mu\mu}$ and $\epsilon^{uA,dA}_{\mu\mu}$, for which confidence intervals had been obtained from oscillation and deep-inelastic scattering measurements [22]. For such ranges, we find a significant presence of NSI in differential and integrated NC (anti)neutrino-nucleon cross sections, particularly for a proton target and at low $Q^{2}$. For this choice of target and kinematics, NSI have a large impact on the polarization of the outgoing nucleons, especially in the transverse direction. We have estimated the uncertainties in the current determination of (flavor-decomposed) axial and strange vector form factors, finding that they leave ample room to study NSI.